OpenStudy (anonymous):

x+2y=8 2x+y=7 solve using substitution

OpenStudy (anonymous):

Okay, so first, we have to get y by itself in one of the equations. So, we'd do:\[-2x + 2x + y = 7 - 2x\] Which will give us:\[y = 7 - 2x\] Then we plug that into the first equation;\[x + 2(7 - 2x) = 8\] Can you try to solve from here?

OpenStudy (anonymous):

x+14-4x=8?

OpenStudy (anonymous):

Yes, that is correct. Now, we add like terms to get:\[-3x + 14 = 8\] Next, we need to get x by itself, so we do: \[-3x + 14 - 14 = 8 - 14\] To get:\[-3x = -6\] And then, we divide -3 from the x by both sides:\[\frac{ -3x }{ -3 } = \frac{ -6 }{ -3 }\] Which would give us what?

OpenStudy (anonymous):

I figured the problem out and got (2,3) maybe you can help me with this one x+3y=2 2x-3y=22

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

Is this substitution or to solve in a different way?

OpenStudy (anonymous):

solve the system using substitution same as last one